OpenStudy (anonymous):

Tonisha leaves Maryville at 7:15 a.m. headed back to college after summer break. Since she is towing a trailer with all of her belongings, she is limited to an average speed of 45 mph. Her friend Sheila leaves Maryville an hour later taking the same route averaging the speed limit of 60 mph. At what time will Sheila pass Tonisha?

OpenStudy (anonymous):

45x = 60 (x - 1) 45x = 60x - 60 15x = 60 x = 4 7 : 15 + 4 hours is 11: 15